OpenStudy (cailindevereux):

Put the function P = 112(0.5)t in the form P = P0ekt. (Round your answers to four decimal places.) P0 = k ≈

OpenStudy (danjs):

It says to turn the exponential with a base 0.5 into one with a base 'e'. \[\large P=112*(0.5)^t\] \[\large P=P _{0}e ^{kt}\]

OpenStudy (danjs):

i would just use a couple data points to compare the two forms... From the given equation with base 0.5 for t=0, P=112*1 = 112 for t=1, P=112*0.5 = 56 Use those values in the base e exponential to get values for Po and k.

OpenStudy (danjs):

Two data points used \[P=P _{0}e^{kt}\] (t,P) = (0, 112) \[\large 112=P _{0}*e^{k*0}~~~~~so~~~~~P _{0}=112\] (t,P)=(1,56) \[P=112e ^{kt}\] \[\large 56=112*e ^{k*1}\]

OpenStudy (danjs):

solving that k=ln(1/2) = -0.693 That is it, \[\large P=112*(0.5)^t\] same as \[\large P=112*e^{-0.693*t}\]

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)